A test has a combination of true/false questions, worth 3 points each, and multiple-choice questions, worth 11 points each. The test has a total of 20 questions and a total of 100 points. How many multiple-choice questions are on the test?
Answer by TimothyLamb(4379)

x = number of true/false questions
y = number of multiple-choice questions
---
x + y = 20
3x + 11y = 100
---
put the system of linear equations into standard form
---
x + y = 20
3x + 11y = 100
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = 15
y = 5
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---